OpenStudy (anonymous):

A charge q moves from point A to point B. The potential energy of the charge at point A is 5.3 × 10-10 joules and at point is B is 3.1 × 10-10 joules. If the kinetic energy of the charge at point A is zero, what is the kinetic energy at point B? Select one of the options below as your answer: A. 8.4 × 10-10 joules B. 5.3 × 10-10 joules C. 3.1 × 10-10 joules D. 2.2 × 10-10 joules E. 1.5 × 10-10 joules

OpenStudy (fifciol):

\[E_k = \Delta E_p=( 5,3-3,1) * 10^{-10} = 2,2*10^{-10} J\]

OpenStudy (anonymous):

SWEET :)

OpenStudy (anonymous):

ty